OpenStudy (anonymous):

Currently, Sam's coach is three times older than Sam. When Sam started soccer 15 years ago, his coach was six times his age. How old is Sam now?

OpenStudy (anonymous):

let Sam's age now be x. and sam's coach age now be y Form 2 eqns y=3x -----(1) y-15= 6(x-15) y-15= 6x-90 y=6x-75 ----(2) substitute (1) into (2) 3x= 6x-75 75=3x x=25 Sam's current age now is 25 years old